OpenStudy (anonymous):

write the equation of the line in point-slope form. (-2,4) and (-5,6). i found the slope, it is 2/-3.

OpenStudy (anonymous):

I am looking for the y-intercept so this is what i have so far:

OpenStudy (anonymous):

y= mx+b 4= 2/-3(-2)+b

OpenStudy (anonymous):

then its 4= -4/-3 +b no idk what to do

OpenStudy (anonymous):

this site helps sometimes try it http://www.wolframalpha.com/

OpenStudy (anonymous):

but if you want an explanation ask someone

OpenStudy (anonymous):

it gave me 4= b+ 4/3 which is 8/3

OpenStudy (anonymous):

now how do i use the slope 2/-3 and b which is 8/3 to make an equation of the line in point-slope form

OpenStudy (ivettef365):

ok you have 4 = -2/3 (-2) + b 4 = 4/3 + b 4 4 12 -4 8 -- - --- = ----- = ------ 1 3 3 3 so b = 8/3

OpenStudy (ivettef365):

now just plug in on formula y = mx + b y = -2/3x + 8/3

OpenStudy (anonymous):

noo i need point slope formula.

OpenStudy (anonymous):

"2.Determine the equation of the line in point-slope form"

OpenStudy (ivettef365):

y - y1 = m(x - x1) (this formual ?)

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

then you have m = -2/3 and (x,y) could be any known point you have point (-2,4) so it is y - 4 = -2/3(x + 2)

OpenStudy (anonymous):

so.. y-4= -2/3 x then what

OpenStudy (phi):

If they want point - slope form, you do not need to first find y= mx + b (slope-intercept form) instead, find the slope, which you did and then pick one of the points, and use that point in \[ y - y_1= m(x-x_1) \]

OpenStudy (anonymous):

y-4= -2/3 x + -4/3 right?

OpenStudy (anonymous):

ohh so i did an extra step woops

OpenStudy (phi):

Don't distribute. point-slope form is just \[ y - y_1= m(x-x_1) \] leave it that way

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

r u sure

OpenStudy (ivettef365):

is just y - 4 = -2/3(x + 2)

OpenStudy (anonymous):

y - y1 = m(x - x1) do not distribute...just leave it , parenthesis and all

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have a few more questions im gonna put the next one

OpenStudy (phi):

point-slope form is not very useful (for plotting for example) but it is easier to write down (less algebra).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my next one is Determine the equation of the line, in slope-intercept form

OpenStudy (anonymous):

the points are:

OpenStudy (anonymous):

(-5,6) and (-8,8)

OpenStudy (phi):

this one is y = mx + b which you have done already first find the slope, then solve for b

OpenStudy (anonymous):

for slope i got 2/-3

OpenStudy (anonymous):

so i plug in a point, and the slope, to solve for b?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok ill use (-5,6)

OpenStudy (anonymous):

ok i got 6= -10/-3 +b

OpenStudy (anonymous):

so that's like the same as 6= 10/3 +b right

OpenStudy (phi):

yes. now add -10/3 to both sides

OpenStudy (anonymous):

10/3 u mean

OpenStudy (phi):

6= 10/3 +b add -10/3 to both sides

OpenStudy (phi):

because you want 10/3 - 10/3 = 0 so you get 0+b on the right side

OpenStudy (anonymous):

wait wait ur confusing me, look what ur writing. first u write 10/3 then u write -10/3.. which is it?

OpenStudy (phi):

here is what you did: ok i got 6= -10/-3 +b so that's like the same as 6= 10/3 +b right yes, that is correct: 6 = 10/3 + b now to solve for b, add -10/3 to both sides

OpenStudy (anonymous):

see u did it again

OpenStudy (anonymous):

u wrote "now to solve for b, add -10/3 to both sides" do u see the problem here?

OpenStudy (anonymous):

but just before you wrote 6= 10/3 +b

OpenStudy (anonymous):

y would i add -10/3 to both sides if we just made the fraction simpler by taking away the negatives?

OpenStudy (anonymous):

do i add or subtract

OpenStudy (anonymous):

if would have been less confusing if u said subtract 10/3 from both sides

OpenStudy (anonymous):

so now whats 6 - 10/3 ?

OpenStudy (phi):

you agree we started with y = - 2/3 x + b you picked point (-5,6) and got 6 = - 2/3 * -5 + b which becomes 6 = 10/3 + b so far, so good. now add -10/3 to both sides. Can you do that ?

OpenStudy (phi):

to do fractions, you need a common denominator. in this case, 3 to get a denominator of 3 for 6, multiply 6 times 3/3 you get 18/3 now you can do the fractions

OpenStudy (anonymous):

so b is 18/3 ??

OpenStudy (phi):

you should get b= 6 - 10/3 change 6 to 18/3 b= 18/3 - 10/3 to add or subtract fractions with the same bottom, add or subtract the tops 18 -10 = 8 and put that number over the common denominator

OpenStudy (anonymous):

so 8/3

OpenStudy (phi):

yes

OpenStudy (phi):

It looks like all the points are ending up on the same line y= -2/3 x + 8/3 looks familiar

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

because they are. i made it that way

OpenStudy (anonymous):

we had to pick three points and make them be in the same line as (1,2) the "launch area" for a project.

OpenStudy (phi):

and then you have to find the different ways to write an equation of a line: standard, point-slope, point-interecept ?

OpenStudy (anonymous):

yes